### v2

@@ 3 Motion Synthesis with Space-time Features
The above architecture is used both in the registered setting [30] and the newly introduced unregistered one, but the training process, losses and evaluation metrics need to be adapted to each case. In Section 3.1, we describe the two settings along with the notations, in Section 3.4 we detail the two strategies we used for training ScanTalk, and in Section 3.5 we illustrate the evaluation metrics. The specifics on the encoding and decoding operations are detailed, respectively, in Section 3.2, and Section 3.3.
@@ 3.1 Protocols and Notation
ScanTalk was designed with the goal of animating meshes in arbitrary topologies. In our first proposal [30], any mesh, seen or unseen during training, could be animated at inference time. Yet, during training, meshes within each training sequence had a fixed and consistent mesh structure (registered setting): the sequences are said to be homogeneous. This allowed us to rely on losses that fully exploit the known vertex-wise correspondence. In the new, unregistered setting, this constraint is dropped. Below we formalize the two protocols.
@@
In this setting, meshes in each sequence of the training set $L$ can be of arbitrary topology, that is, given a generic sequence $M_{i}^{gt}=(m_{i}^{0},\dots,m_{i}^{T_{i}-1})$ each mesh $m_{i}^{j},\;j=0,\cdots,T_{i}-1$ can have its own topology, i.e., $m_{i}^{j}\in\mathbb{R}^{V_{i}^{j}\times 3}$, $m_{i}^{k}\in\mathbb{R}^{V_{i}^{k}\times 3}$ with $V_{i}^{j}\neq V_{i}^{k}$ in general. Thus, we need different loss formulations that do not rely on vertex-wise correspondence within the training sequence (Section 3.4.2). Similar to the registered case, both during training and inference, the topology of the meshes in the generated sequence $\widehat{M}_{i}$ is fixed, matching that of the neutral face $m_{i}^{n}$. However, this time it can differ from that of the ground-truth sequence $M_{i}^{gt}$.
@@ 3.2.1 Face Mesh Encoder
Several approaches are available for encoding face meshes, but traditional graph convolution-based models, like [19, 25], encounter limitations with varying graph structures. Thus, we employ DiffusionNet [46], a discretization-agnostic encoder that has proven effective for encoding face meshes as in [40]. DiffusionNet integrates multi-layer perceptrons (MLPs), learned diffusion, and spatial gradient features, offering a straightforward yet robust architecture for surface learning tasks.
@@ 3.2.2 Audio Encoder
Following [20, 47], the speech is encoded using a pre-trained HuBERT [21] model followed by a Linear layer, with which we obtain a per-frame audio representation:
@@
To ensure temporal coherence in the speech representation, following the methodology outlined in [29], we concatenate the SpeechEncoder with a Multilayer Bidirectional-LSTM, projecting the speech signal into a temporal latent representation:
@@ 3.2.3 Space-time feature field
Properly combining geometric and audio features is of utmost importance, as it will form the base feature space for decoding the final talking head sequence. The main challenge is how to combine these features such that: (i) the process can be done regardless of the mesh topology, and (ii) audio features guide the motion synthesis irrespective of the underlying shape. To do so, we chose to concatenate per-vertex descriptors $f_{i}^{n}$ extracted from the neutral face with temporal latent vectors $\hat{a_{i}}^{j}$ extracted from the Bidirectional-LSTM. Specifically, for each frame (or time instant) $j$, the same audio features $\hat{a_{i}}^{j}$ are replicated and concatenated for each vertex $k$. Conversely, the same geometric features of the neutral face $f_{i}^{n}$ are used for each frame $j$. More formally, this is defined as:
$(F_{i}^{j})_{k}=(f_{i}^{n})_{k}\oplus\hat{a_{i}}^{j}\in\mathbb{R}^{h*2}\qquad\forall\;k=0,\dots,V_{i-1};\;\;\;\forall\;j=0,\dots,T_{i-1}.$ (5)
where $k$ is a mesh vertex, $j$ is a frame in the sequence, and $i$ a sample in the dataset. This structure can be applied to neutral faces with arbitrary topology. Furthermore, since the geometric features are the same for each frame, the motion is guided by the audio features, With this concatenation, we obtain a combined latent $F_{i}^{j}\in\mathbb{R}^{V_{i}\times h*2}$ that embeds both audio-related and geometry-related features. This allows us to propagate the dimensionality of the mesh, $V_{i}$, to the decoder, enabling the desired animation regardless of the mesh structure. This simple concatenation proved to be efficient enough compared to more elaborated strategy such as using an attention mechanism as we show further in Table 3.
@@ 3.3 Decoding the talking sequence
Here, $\widehat{m}_{i}^{j}$ represents the $j$-th frame of the predicted sequence. The entire generated sequence is defined by $\widehat{M}_{i}\in\mathbb{R}^{T_{i}\times V_{i}\times 3}$. To clarify the process, suppose we want to animate a 3D face with $V_{i}$ vertices, $V_{i}$ is arbitrary. The DiffusionNet encoder $DN_{e}$ processes each vertex of $m_{i}^{n}$ along with precomputed features $P_{i}^{n}$, producing a descriptor array of shape $[V_{i},h]$. These descriptors are then concatenated with the descriptors $v_{i}$ from the audio processing pipeline. Each audio feature has the same size $h$ of the geometric features, and are replicated for each vertex, resulting in a combined feature set of shape $[V_{i},h*2]$ per audio frame. This combined feature set, along with the precomputed operators $P_{i}^{n}$ from the input mesh, is fed into the DiffusionNet decoder. The decoder predicts the displacements from the input mesh, yielding an output array of shape $[V_{i},3]$ per audio frame.
We utilize ScanTalk to predict the deformation of the neutral face $m_{i}^{n}$ rather than predicting the actual face. This decision aligns with previous works [14, 20, 35, 47, 50], offering advantages in terms of training efficiency and resulting animation. Predicting face deformation simplifies the learning process by focusing solely on speech-related motion, rather than the entire face reconstruction. The decoder thus learns to predict a per-vertex displacement field from a time-dependent per-vertex descriptors field.
@@ 3.4 Training strategy
The architecture of ScanTalk allows both registered and fully unregistered training. This is different with respect to our previous version in [30], where meshes within a training sequence needed to adhere to a consistent topology (Section 3.1). However, due to its design, the generated deformation field maintains the same topology as the neutral face given as input for animation (which can be arbitrary). Expanding on the discussion in Section 3.1, this distinct feature ensures that even in the unregistered setting the model outputs a sequence of meshes that preserve the topology of the input neutral mesh, while being trainable with per-frame unregistered sequences. Thus, if a sequence is composed of $N$ meshes, each with a different topology, the generated sequence will maintain that of the neutral face. Implicitly, it performs a sort of dense registration, which is a valuable byproduct. This consistency is achieved through the DiffusionNet Decoder, which requires knowledge of precomputed features $P_{i}^{n}$ and the number of points in the animated neutral face. In the following, we report the training details for the two settings, i.e., registered and unregistered.
@@ 3.4.1 Registered Setting
In the registered setting, the ground-truth meshes within each sequence must share a common topology. Yet, different sequences can have meshes of different structures. Despite this requirement, it proves to be non-restrictive in practice as at inference time, any mesh can be animated. The topology can also differ from those seen during training. Moreover, the advantage is that one can exploit the additional knowledge of the mesh structure to impose specific losses or sophisticated constraints.
The model predicts per-vertex displacements of the neutral ground-truth mesh. In a registered setting, the ground truth sequence is registered. As a result, the predicted and ground-truth sequences share the same topology allowing for the definition of specific losses during training. We employed four different losses:
@@ 3.4.2 Unregistered Setting
$d_{CD}(m_{i}^{j},\widehat{m}_{i}^{j})=\frac{1}{\widehat{V}_{i}}\sum_{l}\min_{k}\|(m_{i}^{j})_{k}-(\widehat{m}_{i}^{j})_{l}\|_{2}^{2}+\frac{1}{V_{i}}\sum_{k}\min_{l}\|(m_{i}^{j})_{k}-(\widehat{m}_{i}^{j})_{l}\|_{2}^{2}.$ (13)
@@
Whereas other more sophisticated options exist to compute the discrepancy between generic meshes, in our experiments this loss proved efficient enough in comparison with other losses such as those based on varifolds or optimal transport [3, 9]. The latter are expensive to compute, even for static objects. For speech-driven dynamics, this is even more exacerbated, making them unsuitable.
@@ 3.5 Proposed Evaluation Metrics
In the realm of speech-driven 3D talking head animation, quantitatively assessing the quality of animated faces poses significant challenges. To define the proposed set of metrics, we mainly consider the case where ground-truth and generated meshes share the same topology regardless of the training scheme, for two reasons. First, we are interested in assessing the quality of the animations. Unregistered meshes require first to establish point correspondence, which is an error-prone process. Fully registered meshes allow for a more accurate error evaluation being the point correspondence known. Second, recent research showed how standard measures for evaluating geometric error for unregistered 3D faces can be biased and uninformative [44]. Nonetheless, to fairly evaluate the performance of ScanTalk when used in the fully unregistered setting, we also present results using metrics that do not assume a known topology.
@@ 3.5.1 Registered setting
The most widely used metric for evaluating the accuracy of 3D Talking Head methods is the Lip-Vertex Error (LVE), which calculates the maximum quadratic error between the mouth vertices of the ground truth and the prediction. However, focusing on maximum error fails to capture the comprehensive quality of lip synchronization, as it may be skewed by a single inaccurate vertex, while overall lip-sync quality remains high. As a result, LVE is insufficient for a thorough evaluation of the generated faces and often produces rankings that do not align with qualitative assessments from visual inspections. To address this, we start with the assumption that evaluating lip-sync quality requires assessing the accuracy of lip movements. In the registered setting for training and evaluation, we select six lip vertices (three per lip) for each topology (VOCAset, BIWI, Multiface) and measure the discrepancy between the dynamics of these vertices in the ground truth and the prediction. Specifically, we consider the 3D motion trajectory of these six points throughout the sequences (both generated and real). These can be interpreted as 3D signals, with each signal element corresponding to the $x$, $y$ and $z$ coordinates of the lip vertex in the respective face mesh of the sequence. This approach yields a collection of 3D signals that effectively capture lip movements, allowing for the application of time-series evaluation metrics. To assess the quality of lip synchronization, we propose leveraging two metrics commonly used for measuring distances between polygonal curves: the Discrete Fréchet Distance and Dynamic Time Warping, similar to [50]. Furthermore, we advocate for assessing both the magnitude and angular deviations across subsequent frames. We thus define the Mean Squared Displacement Error ($\delta_{M}$), and Cosine Displacement Error $(\delta_{Cd})$.
Given two sequences of points $P=\{p_{1},p_{2},\ldots,p_{n}\}$ and $Q=\{q_{1},q_{2},\ldots,q_{n}\}$, where $p_{i}$ and $q_{j}$ are points in $\mathbb{R}^{3}$, the following metrics are defined:
Dynamic Time Warping (DTW) [27] ×$\times 10^{-2}$ $(mm)$ measures the similarity between two sequences by allowing for non-linear alignments and different progression rates:
@@
Discrete Fréchet Distance (DFD) [13] ×$\times 10^{-3}$ $(mm)$ measures the similarity between two sequences of points by considering their order. It is defined as:
@@
where $\sigma$ ranges over all possible monotonic bijections between the indices of $P$ and $Q$. A monotonic bijection $\sigma$ is a sequence of index pairs $(i,j)$ such that $1\leq i\leq n$, $1\leq j\leq m$, and $i$ and $j$ are non-decreasing.
Mean Squared Displacement Error $\mathbf{\delta_{M}}$$\times 10^{-6}$ $(mm)$ quantifies the discrepancy between two sequences of 3D points by measuring the magnitude of errors between the displacement vectors of two consecutive frames. It is defined as:
[…]
@@
Cosine Displacement Error $\mathbf{\delta_{Cd}}$ $(rad)$ evaluates the angular similarity between two sequences of 3D points by measuring the cosine distance between the displacement vectors of two consecutive frames. It is defined as:
[…]
@@ 3.5.2 Unregistered setting
In this setting, the generated and ground-truth meshes differ in topology. In fact, all the generated meshes adhere to the input neutral mesh topology, which can differ from the ground-truth. In order to evaluate this scenario, since we do not have any known point-wise correspondence, we use three different shape distances: the chamfer metric previously introduced, the Hausdorff distance (HD) and a kernel metric $\mathcal{L}^{K}$ on varifold representations of meshes introduced in [6, 23]. For two meshes $X$ and $\hat{X}$, it is essentially computed by blending areas $(a_{f})$, centers $(c_{f})$ and normals $(\vec{n}_{f})$ of triangles faces with kernels $k_{p}$ and $k_{n}$:

### v3

@@ 3 Motion Synthesis with Space-time Features
The above architecture is used both in the registered setting [nocentini2024scantalk3dtalkingheads] and the newly introduced unregistered one, but the training process, losses and evaluation metrics need to be adapted to each case. In Section 3.1, we describe the two settings along with the notations, in Section 3.4 we detail the two strategies we used for training ScanTalk, and in Section 3.5 we illustrate the evaluation metrics. The specifics on the encoding and decoding operations are detailed, respectively, in Section 3.2, and Section 3.3.
@@ 3.1 Protocols and Notation
ScanTalk was designed with the goal of animating meshes in arbitrary topologies. In our first proposal [nocentini2024scantalk3dtalkingheads], any mesh, seen or unseen during training, could be animated at inference time. Yet, during training, meshes within each training sequence had a fixed and consistent mesh structure (registered setting): the sequences are said to be homogeneous. This allowed us to rely on losses that fully exploit the known vertex-wise correspondence. In the new, unregistered setting, this constraint is dropped.
@@
In this setting, meshes in each sequence of the training set $L$ can be of arbitrary topology, that is, given a generic sequence $M_{i}^{gt}=(m_{i}^{0},\dots,m_{i}^{T_{i}-1})$ each mesh $m_{i}^{j},\;j=0,\cdots,T_{i}-1$ can have its own topology, i.e., $m_{i}^{j}\in\mathbb{R}^{V_{i}^{j}\times 3}$, $m_{i}^{k}\in\mathbb{R}^{V_{i}^{k}\times 3}$ with $V_{i}^{j}\neq V_{i}^{k}$ in general. Thus, we need different loss formulations that do not rely on vertex-wise correspondence within the training sequence (Section 3.4.2). Similar to the registered case, both during training and inference, the topology of the meshes in the generated sequence $\widehat{M}_{i}$ is fixed, matching that of the neutral face $m_{i}^{n}$. However, this time can differ from that of the ground-truth sequence $M_{i}^{gt}$.
@@ 3.2.1 Face Mesh Encoder
Several approaches are available for encoding face meshes, but traditional graph convolution-based models, like [gong2019spiralnet++, spiralconv_Lim_2019], encounter limitations with varying graph structures. Thus, we employ DiffusionNet [sharp2021diffusion], a discretization-agnostic encoder that has proven effective for encoding face meshes as in [Qin_2023_NFR]. DiffusionNet integrates multi-layer perceptrons (MLPs), learned diffusion, and spatial gradient features, offering a straightforward yet robust architecture for surface learning tasks.
@@ 3.2.2 Audio Encoder
Following [facexhubert, FaceDiffuser_Stan_MIG2023], the speech is encoded using a pre-trained HuBERT [Hubert_audio_encoding] model followed by a Linear layer, with which we obtain a per-frame audio representation:
@@
To ensure temporal coherence in the speech representation, following the methodology outlined in [landmarks_3D_Nocentini_2023], we concatenate the SpeechEncoder with a Multilayer Bidirectional-LSTM, projecting the speech signal into a temporal latent representation:
@@ 3.2.3 Space-time feature field
Properly combining geometric and audio features is of utmost importance, as it will form the base feature space for decoding the final talking head sequence. The main challenge is how to combine these features such that: (i) the process can be done regardless of the mesh topology, and (ii) audio features guide the motion synthesis irrespective of the underlying shape. To do so, we chose to concatenate per-vertex descriptors $f_{i}^{n}$ extracted from the neutral face with temporal latent vectors $\hat{a_{i}}^{j}$ from the Bidirectional-LSTM. Specifically, for each frame (or timestep) $j$, the same audio features $\hat{a_{i}}^{j}$ are replicated and concatenated for each vertex $k$. Conversely, the same geometric features of the neutral face $f_{i}^{n}$ are used for each frame $j$. More formally:
$(F_{i}^{j})_{k}=(f_{i}^{n})_{k}\oplus\hat{a_{i}}^{j}\in\mathbb{R}^{h*2}\qquad\forall\;k=0,\dots,V_{i-1};\;\;\;\forall\;j=0,\dots,T_{i-1},$ (5)
where $k$ is a mesh vertex, $j$ is a frame in the sequence, and $i$ a sample in the dataset. This structure can be applied to neutral faces with arbitrary topology. Furthermore, since the geometric features are the same for each frame, the motion is guided by the audio features, We thus obtain a combined latent $F_{i}^{j}\in\mathbb{R}^{V_{i}\times h*2}$ that embeds both audio-related and geometry-related features. This allows us to propagate the dimensionality of the mesh, $V_{i}$, to the decoder, enabling the desired animation regardless of the mesh structure. A simple concatenation proved effective enough as compared to more elaborated strategies, e.g., attention mechanism. A quantitative analysis of different combination options are reported in the supplementary material.
@@ 3.3 Decoding the talking sequence
Here, $\widehat{m}_{i}^{j}$ represents the $j$-th frame of the predicted sequence. The entire generated sequence is defined by $\widehat{M}_{i}\in\mathbb{R}^{T_{i}\times V_{i}\times 3}$. To clarify the process, suppose we want to animate a 3D face with $V_{i}$ vertices, $V_{i}$ is arbitrary. The DiffusionNet encoder $DN_{e}$ processes each vertex of $m_{i}^{n}$ along with precomputed features $P_{i}^{n}$, producing a descriptor array of shape $V_{i},h]$. These descriptors are then concatenated with the descriptors $v_{i}$ from the audio processing pipeline. Each audio feature has the same size $h$ of the geometric features, and is replicated for each vertex, resulting in a combined feature set of shape $V_{i},h*2]$ per audio frame. This combined feature set, along with the precomputed operators $P_{i}^{n}$ from the input mesh, is fed to the decoder. The decoder outputs a per-vertex displacement field from a time-dependent per-vertex descriptor field, yielding an output array of shape $V_{i},3]$ per audio frame. Predicting a deformation field rather than reconstructing the actual geometry of the animated face offers advantages in terms of training efficiency and resulting animation as it simplifies the learning process by focusing solely on speech-related motion.
@@ 3.4 Training strategy
The architecture of ScanTalk allows both registered and fully unregistered training. However, due to its design, the generated deformation field maintains the same topology as the neutral face given as input for animation (which can be arbitrary). Expanding on the discussion in Section 3.1, this distinct feature ensures that even in the unregistered setting the model outputs a sequence of meshes that preserve the topology of the input neutral mesh, while being trainable with per-frame unregistered sequences. Thus, if a sequence is composed of $N$ meshes, each with a different topology, the generated sequence will maintain that of the neutral face. Implicitly, it performs a sort of dense registration, which is a valuable byproduct. This is achieved thanks to the Decoder, which requires knowledge of the precomputed features $P_{i}^{n}$ and number of vertices of the input mesh, ultimately defining its structure.
@@ 3.4.1 Registered Setting
In the registered setting, the ground-truth meshes within each sequence must share a common topology. Yet, different sequences can have meshes of different structures. Nonetheless, it proved to be non-restrictive in practice as at inference time, any mesh can be animated even if its topology was unseen during training. Furthermore, the advantage is that one can exploit the additional knowledge of the mesh structure to impose specific losses or sophisticated constraints. We employed four different losses:
@@ 3.4.2 Unregistered Setting
$d_{CD}(m_{i}^{j},\widehat{m}_{i}^{j})=\frac{1}{\widehat{V}_{i}}\sum_{l}\min_{k}\|(m_{i}^{j})_{k}-(\widehat{m}_{i}^{j})_{l}\|_{2}^{2}+\frac{1}{V_{i}}\sum_{k}\min_{l}\|(m_{i}^{j})_{k}-(\widehat{m}_{i}^{j})_{l}\|_{2}^{2},$ (13)
@@
Whereas other more sophisticated options exist to compute the discrepancy between generic meshes, in our experiments this loss proved efficient enough in comparison with other losses such as those based on varifolds or optimal transport [besnier_varifold_loss_2023, Croquet_Diff_Reg_OT_2021]. The latter are expensive to compute, even for static objects. For speech-driven dynamics, this is even more exacerbated, making them unsuitable.
@@ 3.5 Proposed Evaluation Metrics
In the realm of speech-driven 3D talking head animation, quantitatively assessing the quality of animated faces poses significant challenges. To define the proposed set of metrics, we mainly consider the registered case. The reason is twofold: first, unregistered meshes require first to establish point correspondence, which is an error-prone process. Fully registered meshes allow for a more accurate error evaluation being the point correspondence known. Second, recent research showed how measuring the geometric error for unregistered 3D faces can be biased and uninformative [sariyanidi2023meta]. Nonetheless, to evaluate the performance of ScanTalk when used in the unregistered setting, we also present results using metrics that do not assume a known topology.
@@ 3.5.1 Registered setting
The most widely used metric for evaluating the accuracy of 3D Talking Head methods is the Lip-Vertex Error (LVE), which calculates the maximum quadratic error between the mouth vertices of the ground truth and the prediction. However, focusing on maximum error fails to capture the comprehensive quality of lip synchronization, as it may be skewed by a single inaccurate vertex, while overall lip-sync quality remains high. As a result, LVE is insufficient for a thorough evaluation of the generated faces and often produces rankings that do not align with qualitative assessments from visual inspections. To address this, we start with the assumption that evaluating lip-sync quality requires assessing the accuracy of lip movements. In the registered setting for training and evaluation, we select six lip vertices (three per lip) for each topology (VOCAset, BIWI, Multiface) and measure the discrepancy between the dynamics of these vertices in the ground truth and the prediction. Specifically, we consider the 3D motion trajectory of these six points throughout the sequences (both generated and real). These can be interpreted as 3D signals, with each signal element corresponding to the $x$, $y$ and $z$ coordinates of the lip vertex in the respective face mesh of the sequence. This approach yields a collection of 3D signals that effectively capture lip movements, allowing for the application of time-series evaluation metrics. To assess the quality of lip synchronization, we propose leveraging two metrics commonly used for measuring distances between polygonal curves: the Discrete Fréchet Distance and Dynamic Time Warping, similar to [Thambiraja_2023_ICCV_imitator]. Furthermore, we advocate for assessing both the magnitude and angular deviations across subsequent frames. We thus define the Mean Squared Displacement Error ($\delta_{M}$), and Cosine Displacement Error $(\delta_{Cd})$. Given two sequences of points $P=\{p_{1},p_{2},\ldots,p_{n}\}$ and $Q=\{q_{1},q_{2},\ldots,q_{n}\}$, where $p_{i}$ and $q_{j}$ are points in $\mathbb{R}^{3}$, the following metrics are defined:
Dynamic Time Warping (DTW) [muller2007dynamic] $\times 10^{-2}$ $(mm)$ measures the similarity between two sequences by allowing for non-linear alignments and different progression rates:
@@
Discrete Fréchet Distance (DFD) [eiter1994computing] $\times 10^{-3}$ $(mm)$ measures the similarity between two sequences of points by considering their order. It is defined as:
@@
where $\sigma$ ranges over all the monotonic bijections between the indices of $P$ and $Q$. A monotonic bijection $\sigma$ is a sequence of index pairs $(i,j)$ such that $1\leq i\leq n$, $1\leq j\leq m$, and $i$ and $j$ are non-decreasing.
Mean Squared Displacement $\mathbf{\delta_{M}}$ $\times 10^{-6}$ $(mm)$ measures the magnitude of errors between the displacement vectors of two consecutive frames. It is defined as:
$\text{$\delta_{M}(P,Q)$}=\frac{1}{n-1}\sum_{i=2}^{n}\left\|(p_{i}-p_{i-1})-(q_{i}-q_{i-1})\right\|^{2},$ (17)
@@
Cosine Displacement $\mathbf{\delta_{Cd}}$ $(rad)$ evaluates the angular similarity between two trajectories by measuring the cosine distance between the displacement vectors of two consecutive frames. It is defined as:
$\text{$\delta_{Cd}(P,Q)$}=\frac{1}{n-1}\sum_{i=2}^{n}\left(1-\frac{(p_{i}-p_{i-1})\cdot(q_{i}-q_{i-1})}{\|p_{i}-p_{i-1}\|\|q_{i}-q_{i-1}\|}\right),$ (18)
@@ 3.5.2 Unregistered setting
In this setting, the generated and ground-truth meshes differ in topology. In fact, all the generated meshes adhere to the input neutral mesh topology, which can differ from the ground-truth. In order to evaluate this scenario, since we do not have any known point-wise correspondence, we use three different shape distances: the Chamfer metric previously introduced, the Hausdorff distance (HD) and a kernel metric $\mathcal{L}^{K}$ on varifold representations of meshes introduced in [Charon_varifold_2013, kaltenmark_varifolds]. For two meshes $X$ and $\hat{X}$, it is essentially computed by blending areas $(a_{f})$, centers $(c_{f})$ and normals $(\vec{n}_{f})$ of triangles faces with kernels $k_{p}$ and $k_{n}$: